# Scalable Generalized Bayesian Online Neural Network Training for Sequential Decision Making

### v1

@@ Abstract
We introduce scalable algorithms for online learning and generalized Bayesian inference of neural network parameters, designed for sequential decision making tasks. Our methods combine the strengths of frequentist and Bayesian filtering, which include fast low-rank updates via a block-diagonal approximation of the parameter error covariance, and a well-defined posterior predictive distribution that we use for decision making. More precisely, our main method updates a low-rank error covariance for the hidden layers parameters, and a full-rank error covariance for the final layer parameters. Although this characterizes an improper posterior, we show that the resulting posterior predictive distribution is well-defined. Our methods update all network parameters online, with no need for replay buffers or offline retraining. We show, empirically, that our methods achieve a competitive tradeoff between speed and accuracy on (non-stationary) contextual bandit problems and Bayesian optimization problems.
@@ 1 Introduction
In sequential decision-making problems, such as Bayesian optimization and contextual bandits, uncertainty quantification and efficient belief updates are essential for balancing exploration and exploitation [15]. A standard approach to this is to use Bayesian neural networks (BNNs) [47]. However, exact posterior inference in BNNs, through methods such as Hamiltonian Monte Carlo (HMC), is computationally expensive, limiting its usage in time-sensitive applications like recommender systems, adaptive control, and financial forecasting [56, 43, 45, 8, 62].
A popular alternative to HMC is the use of variational Bayes methods [57, 6, 31]. Most work uses a Gaussian approximation to the posterior. However, this requires $O({D_{\bm{\theta}}}^{2})$ space to store the covariance (or precision matrix), and $O({D_{\bm{\theta}}}^{3})$ time to update it at each step, where ${D_{\bm{\theta}}}$ is the number of parameters in the network. One solution is to approximate the posterior precision matrix with a diagonal-plus-low-rank structure (see, e.g., the LoFi method [10]), which takes $O({D_{\bm{\theta}}}\,d^{2})$ time to update, where $d$ is the low-rank value. Another approach is to use last-layer Bayesian methods, that use a full covariance Gaussian over the final layer weights, but use a point estimate for the hidden layer weights (see e.g., the VBLL method [7]). These methods take $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}})$ time per update, where ${D_{{\boldsymbol{\ell}}}}$ is the number of parameters in the final layer and ${D_{{\bm{h}}}}$ is the number of parameters in the hidden layers. However, they do not capture uncertainty in the hidden layer parameters (which are usually the majority of the network). See Section 6 for further discussion of related work.
In this work, we propose a methodology that combines recursive frequentist inference, for efficient full-rank/low-rank parameter updates, with a generalized Bayesian posterior predictive that we use for sequential decision making. The frequentist part finds estimates of the parameters and quantifies the error variance of these estimates, which is simpler than characterizing densities. More precisely, we do not need to ensure that the updates we make yield a proper target density. However, following the principles of the Martingale posterior [22], we can still construct a valid posterior predictive distribution without invoking the prior-likelihood-posterior update. In summary, our probabilistic method to train neural networks enjoys these properties: (i) efficient closed-form updates (Monte Carlo sampling is not required); (ii) fully online (it avoids replay buffers, which can be problematic in non-stationary problems); (iii) delivers closed-form uncertainty-aware predictions through the posterior predictive.
@@
We evaluate the performance of our methods across various sequential decision-making tasks. In non-stationary neural contextual bandits, our method achieves the highest performance with lowest computational cost compared to baselines. In stationary settings such as Bayesian optimization, our method matches the performance of replay-buffer approaches while achieving a range of Pareto-efficient tradeoffs between runtime and performance. Our code is implemented in Jax and it is publicly available at github.com/gerdm/low-rank-last-layer-inference
@@ 2 Problem statement
Consider a sequence of output measurements ${\bm{y}}_{1:T}=({\bm{y}}_{1},\ldots,{\bm{y}}_{T})$, with ${\bm{y}}_{t}\in^{D_{\bm{y}}}$, which are generated conditional on some input features ${\bm{x}}_{1:T}=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$, with ${\bm{x}}_{t}\in^{{D_{\bm{x}}}}$, for $t\in\{1,\dots,T\}$ and $T\in\mathbb{N}$. Let $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ be a datapoint at time $t$ and $\mathcal{D}_{1:T}=(\mathcal{D}_{1},\ldots,\mathcal{D}_{T})$ the dataset at time $T$. We assume that ${\bm{y}}_{t}$, conditional on ${\bm{x}}_{t}$, is generated from a distribution that is characterized with unknown parameters ${\bm{\theta}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$, i.e., ${\bm{y}}_{t}\sim p({f}({\bm{\theta}}_{t},{\bm{x}}_{t}))$, for some observation model $p$ with mean modelled by ${f}$.
[…]
Our goal is to efficiently compute best linear unbiased predictions (BLUPs) under a linearized model of ${f}$, along with an approximate block-diagonal error variance-covariance (EVC) structure that captures uncertainty separately for the hidden and last layer parameters. While we do not model or approximate the full Bayesian posterior $p({\bm{\theta}}_{t}\mid\mathcal{D}_{1:t})$, we treat the BLUP parameter estimates and their EVCs as defining degenerate Gaussian distributions [54], which we use to construct a proper posterior predictive $p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t})$ that supports probabilistic, sample-based sequential decision making. In this sense, we combine frequentist recursive parameter estimation with Bayesian decision making via the posterior predictive.
@@ 3.1 Tools for sequential decision making
A key challenge in sequential decision making is the exploration-exploitation tradeoff, in which an agent must balance the use of its current knowledge of the environment (exploitation) with acquiring new information to improve future decisions (exploration) [4]. For example consider the case of a $K$-armed contextual bandit. Let the input (context) be ${\bm{x}}_{t}$, and the predicted mean reward for arm $a$ be ${f}({\bm{\theta}}_{t},{\bm{x}}_{t},a)$, where $a\in\{1,\ldots,K\}$. A simple approach — which we use as a baseline — is the $\epsilon$-greedy exploration, in which the next action is $a_{t+1}=\operatorname*{arg\,max}_{a}{f}(\hat{{\bm{\theta}}}_{t},{\bm{x}}_{t},a)$ with probability $1-\epsilon$, and uniformly at random with probability $\epsilon$, where $\epsilon\in(0,1)$ and $\hat{{\bm{\theta}}}_{t}$ is a point estimate.
A more sophisticated approach is known as Thompson sampling (TS) [52], where an agent samples the parameters from the posterior distribution (or approximate posterior [50]) and then makes a greedy decision based on a plugin estimate of the reward given the sampled parameter. This approach is studied in the context of Bayesian neural bandits in [51]. In this work, however, we focus on uncertainty in predictions rather than uncertainty in parameters. For example, for the case of the $K$-armed contextual bandit, our approach works as follows: at each timestep $t$, estimates of the reward are sampled from the posterior predictive $\hat{y}_{a}\sim p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},a,\mathcal{D}_{1:t})$, after which an action is selected as $a_{t+1}=\operatorname*{arg\,max}_{a\in\{1,\ldots,K\}}\hat{y}_{a}$.222 An alternative design is for ${f}$ to output a reward for each action, i.e., ${f}:\mathbb{R}^{D_{\bm{\theta}}}\times\mathbb{R}^{K\,{D_{\bm{x}}}}\to\mathbb{R}^{K}$. In this case, ${\bm{x}}_{t}\in\mathbb{R}^{K,{D_{\bm{x}}}}$ represents the concatenated features for all arms. Both configurations are explored in our experiments.
This predictive-based approach to TS is also considered in the context of Bayesian optimization in [7]. It is particularly relevant in our setting because our parameter estimates resemble degenerate multivariate Gaussians, due to the positive semi-definiteness of the covariance matrices [54]. Despite this, we show that the posterior predictive distribution is a valid density, (cf. [22]).
3.2 Extended Kalman filtering for online learning
We model observations as ${\bm{y}}_{t}={f}({\boldsymbol{\ell}}_{t-1},{\bm{h}}_{t-1},{\bm{x}}_{t})+{\bm{e}}_{t}$, where ${f}$ is a neural network and ${\bm{e}}_{t}$ is zero-mean noise with covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$; see Appendix A for notation. Next, assume that the latent parameter vector ${\bm{\theta}}_{t}=[{\boldsymbol{\ell}}_{t}^{\intercal},{\bm{h}}_{t}^{\intercal}]^{\intercal}$ evolves as ${\bm{\theta}}_{t}={\bm{\theta}}_{t-1}+{\bm{u}}_{t}$ where ${\bm{u}}_{t}$ is a zero-mean dynamics noise process with variance $\mathbf{Q}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$. Following the frequentist interpretation of the Kalman filter [28, 21], we aim to recursively compute the best linear unbiased predictor (BLUP) ${\bm{\theta}}_{t|k}=\mathbf{A}_{t|k}{\bm{y}}_{1:k}$ and its associated error variance-covariance (EVC) matrix $\boldsymbol{\Sigma}_{t|k}={\rm Cov}({\bm{\theta}}_{t}-{\bm{\theta}}_{t|k})$, for $k=t$ and $k=t-1$, under a local linearization of the model. Here, $\mathbf{A}_{t|k}\in{\mathcal{M}_{{D_{\bm{\theta}}}\times(k\,{D_{\bm{y}}})}(\mathbb{R})}$ is the best linear mapping (in $L^{2}$) from the collection of the first $k$ observations to the space of parameters at time $t$.
To obtain closed-form solutions for the BLUP and EVC at time $t$, we linearize the observation model ${f}$ around ${\bm{\theta}}_{t-1|t-1}=({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1})$ to obtain
${\bm{y}}_{t}\approx{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{L}}_{t}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t-1|t-1})+\tilde{\mathbf{H}}_{t}({\bm{h}}_{t}-{\bm{h}}_{t-1|t-1})+{\bm{e}}_{t},$ (1)
where $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$ and $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$. It then follows that ${\bm{\theta}}_{t|t}$ and $\boldsymbol{\Sigma}_{t|t}$ can be written as a function of ${\bm{\theta}}_{t|t-1}$ and $\boldsymbol{\Sigma}_{t|t-1}$, which can be written as a function of ${\bm{\theta}}_{t-1|t-1}$ and $\boldsymbol{\Sigma}_{t-1|t-1}$ and so on. As a consequence, these updates enable efficient sequential (single-pass) online learning. This procedure is known as the extended Kalman filter (EKF).

### v2

# Martingale Posterior Neural Networks for Fast Sequential Decision Making

@@ Abstract
We introduce scalable algorithms for online learning of neural network parameters and Bayesian sequential decision making. Unlike classical Bayesian neural networks, which induce predictive uncertainty through a posterior over model parameters, our methods adopt a predictive-first perspective based on martingale posteriors. In particular, we work directly with the one-step-ahead posterior predictive, which we parameterize with a neural network and update sequentially with incoming observations. This decouples Bayesian decision-making from parameter-space inference: we sample from the posterior predictive for decision making, and update the parameters of the posterior predictive via fast, frequentist Kalman-filter-like recursions. Our algorithms operate in a fully online, replay-free setting, providing principled uncertainty quantification without costly posterior sampling. Empirically, they achieve competitive performance–speed trade-offs in non-stationary contextual bandits and Bayesian optimization, offering 10–100 times faster inference than classical Thompson sampling while maintaining comparable or superior decision performance.
@@ 1 Introduction
In various sequential decision-making problems, such as Bayesian optimization and contextual bandits, uncertainty quantification and efficient belief updates are essential for balancing exploration and exploitation. A prominent Bayesian approach is Thompson sampling (TS) [52], which samples from the posterior over model parameters and evaluates the corresponding predictive function (e.g., reward or surrogate) for decision-making.
Various works combine Bayesian uncertainty with the expressiveness of neural networks for sequential decision making [5, 51]. These methods, often called Bayesian neural networks (BNNs), typically approximate the posterior over model parameters (e.g., via variational inference) and then use this posterior for downstream tasks. However, BNNs often rely on misspecified priors and likelihoods [33], which often degrade predictive performance [60] and result in slower inference than non-Bayesian alternatives [35].
These limitations are especially problematic in online settings where updates and decisions must be made quickly, such as in recommender systems, adaptive control, financial forecasting, and large-scale Bayesian optimization [55, 43, 45, 6, 59].
To address these problems, we present a class of algorithms for Bayesian sequential decision making inspired by the predictive-first philosophy of the martingale posterior framework [20, 26]. Our approach operates in a fully online setting: the one-step-ahead posterior predictive is parameterized by a neural network, updated directly from real observations via frequentist Kalman filter-style recursions, and sampled once per step for decision making. This avoids the need for costly posterior inference, while still supporting uncertainty and sampling-based sequential decision making.
The parameters defining the predictive density are updated with a single-pass frequentist Kalman filter-style methods, which resemble online natural gradient steps [48] and require no replay buffer (copies of past observations, which one typically uses for multiple inner-iterations). We explore three structured covariance strategies: LRKF applies low-rank updates to all layers; HiLoFi applies a full-rank update to the last layer and a low-rank update to the hidden layers; LoLoFi applies low-rank updates to both. HiLoFi and LoLoFi are inspired by last-layer Bayesian methods [25], and offer different tradeoffs between computational efficiency and expressiveness.
Decision making is performed by sampling directly from the posterior predictive, resulting in 10 to 100 times faster inference than that of classical TS, even when compared with structured posteriors that use diagonal plus low-rank covariances [8]. Our algorithms scale to million-parameter networks, maintain constant-time updates, and operate in fully online, streaming-compatible settings. In addition, our methodology supports any acquisition strategy, such as classical TS or expected improvement.
In summary, our probabilistic method for training neural networks enjoys the following properties: (i) efficient closed-form updates (no Monte Carlo sampling required); (ii) sample efficiency through low-rank Kalman filter-like updates which do not require the specification of a posterior density; (iii) closed-form uncertainty-aware predictions via the posterior predictive.
@@
We evaluate the performance of our methods across a range of sequential decision-making tasks. In non-stationary neural contextual bandits, our methods achieve the highest performance at the lowest computational cost compared to that of baselines. In stationary settings such as Bayesian optimization, our approach matches the performance of replay-buffer methods while achieving Pareto-efficient tradeoffs between runtime and performance. Our code is available at https://github.com/gerdm/martingale-posterior-neural-networks.
A complete summary of the notation used throughout the paper is in Appendix A.
@@ 2 Problem statement
We consider a sequential-decision-making agent that at time $t$ observes a context-action pair ${\bm{x}}_{t}=({\bm{c}}_{t},{\bm{a}}_{t})$ with context ${\bm{c}}_{t}\in\mathbb{R}^{{D_{c}}}$ (possibly empty) and action ${\bm{a}}_{t}\in{\cal A}\subseteq\mathbb{R}^{D_{a}}$. The environment returns a reward (or observation) according to
${\bm{y}}_{t}\sim p_{\rm env}(\cdot\,|\,{\bm{x}}_{t}),\qquad{\bm{y}}_{t}\in\mathbb{R}^{{D_{\bm{y}}}},$ (1)
which depends only on the context-action pair ${\bm{x}}_{t}$. We use the notation $y_{t}$ when the outcome is scalar (${D_{\bm{y}}}=1$) and ${\bm{y}}_{t}$ when it is vector-valued (${D_{\bm{y}}}>1$).
This formulation covers, e.g., Bayesian optimization (${\bm{c}}_{t}$ empty and ${\bm{a}}_{t}$ the query location), multi-armed bandits (${\bm{c}}_{t}$ empty and ${\bm{a}}_{t}$ the chosen arm), and contextual bandits (${\bm{c}}_{t}$ non-empty and ${\bm{a}}_{t}$ the chosen arm). It also serves as a proxy representation for multi-armed bandits by letting the components of ${\bm{y}}_{t}$ encode per-arm signals or auxiliary outcomes.
We write $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ for a datapoint and $\mathcal{D}_{1:t}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{t}\}$ for the dataset. Here, we assume that $p_{\rm env}$ is unknown, but we approximate it through the observation model
${\bm{y}}_{t}=f({\bm{\theta}}_{t},{\bm{x}}_{t})+{\bm{e}}_{t},$ (2)
with ${\bm{e}}_{t}$ a zero-mean random vector with covariance matrix $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$, $f$ a neural network, and ${\bm{\theta}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ unknown time-varying model parameters.
The agent maintains a belief state ${\bm{b}}_{t}$ that summarizes its knowledge of the environment after observing $\mathcal{D}_{1:t}$. This belief defines a parametric approximation to the one-step-ahead posterior predictive
$p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1})\approx p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t}).$ (3)
Our objective is to maintain ${\bm{b}}_{t}$ via efficient, single-pass frequentist updates and to use the predictive $p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\mid{\bm{x}}_{t+1})$ for probabilistic, sample-based sequential decision making. In this sense, we combine frequentist recursive parameter estimation with Bayesian decision making via the posterior predictive.
@@ 3.1 Tools for sequential decision making
A central challenge in sequential decision making is the exploration-exploitation tradeoff, where an agent balances whether to use its current knowledge of the environment (exploitation) or to acquire new information to improve future decisions (exploration) [3].
For example, when solving Bayesian neural contextual bandit problems [51] with the classical Thompson sampling (TS) algorithm [52], an agent samples a parameter vector from the posterior distribution (or an approximation [50]), evaluates the reward function under this sampled parameter for each action, and selects the action with the highest sampled reward.
Here, we propose a novel approach that avoids sampling from (high-dimensional) parameter posteriors often found in Bayesian neural networks. Instead, we work directly with the posterior predictive distribution defined by the current belief state ${\bm{b}}_{t}$. For each candidate action, we sample a possible outcome from (3) and then choose the action with the highest sampled reward. The belief ${\bm{b}}_{t}$ is subsequently updated via frequentist recursions (Section 3.2).
Table 1 highlights the key differences between our predictive sampling approach and classical TS. The essential distinction is whether uncertainty is represented in parameter space (TS) or directly in outcome space (our approach).
Step Predictive sampling (Our approach) Classical TS
Sample Posterior predictive at each action Posterior over model parameters
Evaluate N/A Sampled parameters on reward function
Select action Argmax over sampled rewards Argmax over sampled rewards
Get reward $y_{t+1}\sim p_{\rm env}(\cdot\mid{\bm{x}}_{t+1})$ $y_{t+1}\sim p_{\rm env}(\cdot\mid{\bm{x}}_{t+1})$
Update belief Frequentist update (e.g., Algorithm 2) Bayesian posterior update over parameters
Table 1: Comparison between Bayesian predictive sampling (our approach) and classical Thompson sampling. Differences highlighted in bold.
Algorithm 1 shows our predictive sampling procedure for contextual bandits. At each decision-making step, the agent samples rewards from the predictive distribution for each action, selects the action with the highest sampled reward, observes the reward from the environment, and updates its belief state.
@@
[…]
4:for all ${\bm{a}}\in{\cal A}$ do
5:   $\hat{{\bm{x}}}\leftarrow({\bm{c}}_{t+1},\,{\bm{a}})$
[…]
7:end for
[…]
12:Return $({\bm{b}}_{t+1},y_{t+1})$ // belief and reward
Algorithm 1 Predictive sampling for sequential decision making in contextual bandits.
3.2 Extended Kalman filtering for online learning and sequential decision making
To maintain and update the belief state ${\bm{b}}_{t}$, we adapt ideas from the frequentist perspective of the extended Kalman filter (EKF). Here, ${\bm{b}}_{t}=({\bm{\theta}}_{t|t},\,\boldsymbol{\Sigma}_{t})$, ${\bm{\theta}}_{t|t}=\mathbf{A}_{t}\,{\bm{y}}_{1:t}$, $\mathbf{A}_{t}=\operatorname*{arg\,min}_{\mathbf{A}}\mathbb{E}[\|{\bm{\theta}}_{t}-\mathbf{A}\,{\bm{y}}_{1:t}\|_{2}^{2}]$ is the best linear unbiased predictor (BLUP) of ${\bm{\theta}}_{t}$, and $\boldsymbol{\Sigma}_{t}={\rm Var}({\bm{\theta}}_{t}-{\bm{\theta}}_{t|t})$ is the error-variance-covariance (EVC) matrix of the estimate. The initial belief state ${\bm{b}}_{0}=({\bm{\theta}}_{0|0},\,\boldsymbol{\Sigma}_{0})$ is given.
Next, we partition the neural network parameters into last-layer weights and hidden-layer weights: ${\bm{\theta}}_{t}=({\boldsymbol{\ell}}_{t},{\bm{h}}_{t})$, with ${\boldsymbol{\ell}}_{t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ (last layer) and ${\bm{h}}_{t}\in\mathbb{R}^{D_{{\bm{h}}}}$ (hidden layers) so that ${\bm{\theta}}_{t|t}=({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t})$, and ${f}({\boldsymbol{\ell}}_{t},{\bm{h}}_{t},{\bm{x}}_{t})={\boldsymbol{\ell}}_{t}^{\intercal}\,\phi({\bm{x}}_{t},{\bm{h}}_{t})\in\mathbb{R}$, where $\phi({\bm{x}}_{t},{\bm{h}}_{t})$ is the hidden-layer feature map of the neural network.
To obtain closed-form updates, we linearize the neural network around the previous parameter estimate ${\bm{\theta}}_{t-1|t-1}=({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1})$, and make a random-walk assumption for the model parameters ${\bm{\theta}}_{t}={\bm{\theta}}_{t-1}+{\bm{u}}_{t}$, with ${\rm Var}({\bm{u}}_{t})=\mathbf{Q}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{\theta}}}}(\mathbb{R})}$. This assumption prevents uncertainty collapse, adaptation in non-stationary environments, and numerical stability (see e.g., [46] or Section 2.6.4 in [13]).
With the above assumptions, we obtain the linearized state-space model
${\bm{\theta}}_{t}$ $={\bm{\theta}}_{t-1}+{\bm{u}}_{t},$ (4)
where ${\bm{e}}_{t}$ is a zero-mean random variable with ${\rm Var}[{\bm{e}}_{t}]=\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$, $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})\in{\mathcal{M}_{1\times 2}(\mathbb{R})}$, and $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$ are the Jacobians of the neural network w.r.t. the last and hidden layers respectively.
Updates.
The state-space model defined by (4)-(5) allows for recursive estimation of ${\bm{b}}_{t}$ via Kalman-like updates. See Appendix B for details.
Posterior predictive model.
At each timestep the one-step-ahead posterior predictive induced by the state-space model (4)-(5), given the belief state ${\bm{b}}_{t}$ and under the assumption $p({\bm{e}}_{t})={\cal N}({\bm{e}}_{t}\mid\boldsymbol{0},\,\mathbf{R}_{t})$ is
$p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\mid{\bm{x}}_{t+1})=\mathcal{N}\,\left({\bm{y}}_{t+1}\,\middle|\,{f}({\bm{\theta}}_{t|t},{\bm{x}}_{t+1}),\;\underbrace{\tilde{\mathbf{J}}_{t+1}\,\boldsymbol{\Sigma}_{t}\,\tilde{\mathbf{J}}_{t+1}^{\intercal}}_{\text{epistemic}}+\underbrace{\mathbf{R}_{t}}_{\text{aleatoric}}\right),$ (6)
where $\tilde{\mathbf{J}}_{t}=\begin{bmatrix}\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}_{t}\end{bmatrix}$. Here, the covariance of the posterior predictive decomposes into epistemic uncertainty (from parameter uncertainty under the linearization) and aleatoric uncertainty (from observation noise $\mathbf{R}_{t}$). The condition $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$ ensures that (6) defines a proper Gaussian density.
3.3 Martingale posteriors
Our notion of Bayesian uncertainty is inspired by the martingale posterior framework [20], which interprets Bayesian uncertainty through missing data [26]. In the original setting, one imagines imputing an infinite sequence of future observations; statistics computed on these extended datasets form a martingale, and under mild conditions they converge.
However, in sequential decision making, however, given the context-action pair ${\bm{x}}_{t+1}$, the only missing data, is the next observation ${\bm{y}}_{t+1}$ from the environment. In our case, uncertainty is quantified to guide decisions: ${\bm{y}}_{t+1}$ (or a partial observation of it, in the bandit setting) is always revealed before the next step. Thus, rather than simulating an infinite sequence of future steps, we focus entirely on the one-step-ahead posterior predictive $p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\mid{\bm{x}}_{t+1})$.